The area of a rectangular garden is 84 square meters. The width is 3 meters longer than one-third of the length. Find the length and the width of the garden. Use the formula, area=length*width. Thanks for the help!!
Answer by amarjeeth123(569)

Let the length be 3x.
Then the width is (x+3).
area=length*width
3x(x+3)=84
x(x+3)=28
x^2+3x-28=0
x^2+7x-4x-28=0
x(x+7)-4(x+7)=0
(x-4)(x+7)=0
x=4 is the only positive solution.
The length is 12 metres.
The width is 7 metres.